OpenStudy (anonymous):

MEDAL AND FAN!!!!!!!!! Consider the function f(x) = 4x^2 + 5. Find the value of f(x) when x = 2 and select the correct answer below

OpenStudy (anonymous):

4*2^2+5 4*4+5 16+5 21

OpenStudy (anonymous):

Which of the following expressions represents a function

OpenStudy (anonymous):

x=7 2x^2 +y^2=5 y^2=16-x^2

OpenStudy (anonymous):

Don't forget to medal and fan if you already haven't. Plus mention me if you need more help @hamza123

OpenStudy (anonymous):

A

OpenStudy (anonymous):

which one above is a function

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Consider the function f(x) = 8 − 3X^2. Find the value of f(x) when x = 3 and select the correct answer below

OpenStudy (anonymous):

8-3*3^2=-19

OpenStudy (anonymous):

if u want multiple questions just repost it as a new question so u can medal